Rf=2%, Rm= 10%
An equal weighted portfolio of stock X and Rf yields a return of 10%

A portfolio of stock Y and Rf with a 75% investment in stock Y yields a return of 6.5%
What is the market beta of a portfolio if we invest 40%in X, 20% in Y, 20% in Rf and 20% in the market?
Answer #1

0.5*Rx+0.5*2%=10% =>Rx=18%
0.75*Ry+0.25*2%=6.5% =>Ry=8%

Beta=40%*(18%-2%)/(10%-2%)+20%*(8%-2%)/(10%-2%)+20%*0+20%*1=1.15